### v1

# Dance-to-Music Generation with Encoder-based Textual Inversion of Diffusion Models

Abstract
The harmonious integration of music with dance movements is pivotal in vividly conveying the artistic essence of dance. This alignment also significantly elevates the immersive quality of gaming experiences and animation productions. While there has been remarkable advancement in creating high-fidelity music from textual descriptions, current methodologies mainly concentrate on modulating overarching characteristics such as genre and emotional tone. They often overlook the nuanced management of temporal rhythm, which is indispensable in crafting music for dance, since it intricately aligns the musical beats with the dancers’ movements. Recognizing this gap, we propose an encoder-based textual inversion technique for augmenting text-to-music models with visual control, facilitating personalized music generation. Specifically, we develop dual-path rhythm-genre inversion to effectively integrate the rhythm and genre of a dance motion sequence into the textual space of a text-to-music model. Contrary to the classical textual inversion method, which directly updates text embeddings to reconstruct a single target object, our approach utilizes separate rhythm and genre encoders to obtain text embeddings for two pseudo-words, adapting to the varying rhythms and genres. To achieve a more accurate evaluation, we propose improved evaluation metrics for rhythm alignment. We demonstrate that our approach outperforms state-of-the-art methods across multiple evaluation metrics. Furthermore, our method seamlessly adapts to in-the-wild data and effectively integrates with the inherent text-guided generation capability of the pre-trained model. Samples are available at https://youtu.be/D7XDwtH1YwE.
Figure 1: We propose dual-path rhythm-genre inversion to incorporate rhythm and genre information from a dance motion sequence into two learnable tokens, which are then used to enhance the pre-trained text-to-music models with visual control. Through encoder-based textual inversion, our method offers a plug-and-play solution for text-to-music generation models, enabling seamless integration of visual cues. The word bank comprises learnable pseudo-words (represented as “@” for genre and “*” for rhythm) and descriptive music terms such as instruments, genres, and emotions. By combining these pseudo-words and descriptive music terms, our method allows for the generation of a diverse range of music that is synchronized with the dance.
1 Introducion
If life is a sort of dance, then music is the enchanting melody that accompanies our every graceful movement. Music’s power lies in its ability to synchronize and enhance visual narratives, a trait extensively utilized in gaming and film industries. The advent of short video content has seen a surge in amateur bloggers sharing dance videos on social media, showcasing music’s integral role in this domain. Traditionally, aligning music with video has been labor-intensive, often limiting choreography to align with the music’s attributes. Our study explores using text-to-music generative models to produce music synchronized with dance videos. This approach seeks to offer an artistically and computationally efficient solution, enabling the creation of music that not only complements but also enriches the dance experience.
As AI-generated content continues to advance, there has been a surge in research focusing on dance-to-music generation, aligning with contemporary trends and reflecting the broader evolution of AI applications in creative domains. Some approaches [14, 42, 9, 3, 16] employ symbolic music representation, such as MIDI and REMI, to generate music for dance videos. However, the generated music is often limited to a few classical instruments, such as piano or guitar. Furthermore, these methods often result in incoherent note sequences and simplistic melodies, deviating significantly from the characteristics of real dance music. Other approaches [54, 55, 51, 45] directly use waveforms or mel-spectrograms to generate music for dance videos, resulting in richer musical compositions. However, their audio quality is often limited, and some methods tend to overly emphasize drum beats without capturing melodic elements. Furthermore, these approaches are constrained by the genres prevalent in the dataset, making it challenging to generate diverse music suitable for dance videos. Consequently, they face difficulties when applied to real-world scenarios. Pre-trained text-to-music models have demonstrated their ability to generate high-quality and diverse music. However, most current models only provide control over global attributes of the music and lack the ability to manipulate local properties such as rhythm.
To address the above-mentioned problems, we propose a novel approach for dance-to-music generation based on arbitrary pre-trained text-to-music models. Our method aims to integrate rhythm information into this model while preserving its original text-based generation capability. Given a dance video, the model generates synchronized music and allows for the manipulation of high-level features of music, ensuring a harmonious integration of music and the dance video. As illustrated in Fig. 1, our method enables dance-to-music generation in various genres (e.g., “pop”, “lock”, “house”). Additionally, it provides the ability to control high-level features of music, such as genre and emotion, using textual description. Due to our utilization of motion sequences for rhythm extraction, our method can be extended to various other physical activities (e.g., “jump rope”, “martial arts”, and “artistic gymnastics”).
To achieve this goal, we seek a method to incorporate rhythm information into a pre-trained text-to-music model. This is inspired by Textual Inversion [12], which employs a pseudo-word to represent a specific concept through image reconstruction. We aim to learn a pseudo-word that represents rhythm information of arbitrary dance videos and integrate it into the vocabulary of the text encoder. However, in the text encoder, the text embeddings corresponding to an individual token are fixed, while the rhythms in different videos are variable, making it challenging to create a pseudo-word for each rhythm pattern. Therefore, we propose an encoder-based textual inversion method (inspired by [12]) that extends the text encoder with an encoder branch capable of mapping variable rhythm information to a text embedding. In contrast to prior methods [12, 53, 21, 31] using a fixed text embedding for each pseudo-word, our approach incorporates an encoder branch that enables the text embedding associated with the pseudo-word to dynamically adapt to varying input conditions. As a result, our method enables the generation of music with different rhythms based on different motion sequences in the video.
To narrow the gap between motion sequences and text embeddings, we employ a hybrid approach that combines traditional feature extraction techniques, extracting rhythm sequences from the motion sequences, and projector networks to construct the rhythm encoder. Furthermore, we introduce a genre encoder to provide additional control over music generation.
Our contributions can be summarized as follows:
  • We propose a novel encoder-based textual inversion method that enables a single pseudo-word to represent a class of variable attributes rather than a fixed object.
  • We develop rhythm and genre encoders to achieve dual-path rhythm-genre inversion, converting rhythm and genre information into text embeddings. This enhances text-to-music models by integrating visual control. In dance-to-music generation, it provides flexibility in high-level feature control of music with text.
  • We improve rhythm alignment evaluation metrics to align better with human perception. Experimental results demonstrate that our method achieves optimal performance on multiple evaluation metrics. Furthermore, our method seems to offer seamless adaptation to in-the-wild data.
2 Related Work
Video-aligned audio generation
Video-aligned audio generation has become increasingly popular in recent years. Several approaches have explored the synthesis of sound aligned with videos [22, 33, 10, 43]. Some works utilize symbolic representation for video music generation. Audeo [41] employs a multi-stage approach, while Foley Music [14] combines graph neural networks and Transformers to generate MIDI event sequences for performance videos. RhythmicNet [42] implements a multi-stage approach for dance-to-music generation. Controllable Music Transformer (CMT) [9] designs rule-based rhythmic relationships, allowing control over timing, motion, beat, and music genre for video background music generation. DANCE2MIDI [16] constructs a dataset D2MIDI for multi-instrument MIDI and dance pairing, enabling the generation of coherent music sequences from dance videos. Video2Music [23] and V-MusProd [56] leverage multiple video features and transformer models to generate music that is synchronized with videos. However, symbolic representation-based methods overlook timbre and dynamics, limiting musical expressiveness and variation in the generated music.
Recent studies have investigated the utilization of spectrograms or audio waveforms for video-to-music generation. Dance2Music-GAN (D2M-GAN) [54] is an adversarial multi-modal framework that generates complex music samples conditioned on dance videos. Zhu et al. [55] introduce a novel approach that maximizes mutual information using a conditional discrete contrastive diffusion (CDCD) loss to generate dance music. LORIS (Long-Term Rhythmic Video Soundtracker) [51] employs a latent conditional diffusion probabilistic model and context-aware conditioning encoders for synthesizing long-term conditional waveforms.  Tan et al. [45] combines a UNET-based latent diffusion model and a pre-trained VAE model to generate plausible dance music from 3D motion data and genre labels. However, most of them generate low-quality audio with significant noise, and the generated music tends to be similar, making it challenging to adapt to diverse video scenarios in the wild. In contrast, our approach, based on a pre-trained text-to-music model, is capable of generating diverse and high-quality beat-aligned dance music. V2Meow [40] utilizes a multi-stage autoregressive model to generate visually-aligned music, allowing for high-level feature control with text. However, it requires training from scratch on a large-scale dataset of O(100k), in contrast to our approach, which employs a small-scale dataset of O(1k). By leveraging encoder-based textual inversion, we drive a pre-trained text-to-music model to achieve dance-to-music generation.
Text-to-music generation
Text-to-music generation has made remarkable advances in recent years. Any-to-Any generation such as CoDi [46] and NExT-GPT [48] have the capability to generate across different modalities, including text-to-music generation. Several approaches like Make-an-Audio [20], AudioLDM [28], AudioLDM2 [29], and Tango [15] utilize diffusion models to generate text-guided audio encompassing speech, sounds, and music. However, these methods face limitations in terms of data quality and model scalability, resulting in lower-quality music generation. Some works such as Archisound [38], Riffusion [11], Moûsai [39], and Noise2Music [19] leverage diffusion models to achieve high-quality music generation from textual input. Furthermore, transformer-based frameworks like MusicLM [4], MUSICGEN [8], and MuseCoco [30] encode music as discrete tokens, enabling the generation of high-quality music with rich textual description. However, the above methods mainly control high-level features like genre and emotion in music generation, leaving out control over the rhythm aligned with videos. In contrast, our method utilizes learnable pseudo-words to generate video-aligned music, distinguishing it from previous approaches.
Personalization of generative models
While text-guided content generation has achieved impressive results, relying solely on text is insufficient for precise control over the generated content, especially when targeting specific objectives. Therefore, personalization of generative models has become a recent research focus. Personalization of generative models refers to the generation of content personalized to specific objects based on text-to-content models, leveraging prompt learning or fine-tuning techniques. DreamBooth [37] introduces a class-specific prior preservation loss for personalized image generation. LORA [18] is an efficient fine-tuning method known for achieving remarkable performance in personalized generation. Music ControlNet [49] enhances text-to-music models with composable, precise, and time-varying musical controls, without incorporating visual controls. Gal et al. [12] introduce a textual inversion (TI) method that progressively updates the embeddings of text placeholders corresponding to specific object’s visual features within a pre-trained text encoder. Inspried by TI, many variants [13, 21, 26, 47, 31, 53, 52, 5] achieve high-quality and more controllable personalized image generation. Plitsis et al. [32] directly investigate DreamBooth and TI in text-to-music generation, but this work is limited to editing audio with a reference audio. In dance-to-music generation, the incorporation of rhythm information from dance videos is necessary. Thus, as a new task of TI, we propose an encoder-based textual inversion approach that facilitates personalized music generation, allowing for the generation of music that is aligned with specific rhythmic video.
3 Method
Figure 2: We employ a latent diffusion model as the generative backbone and propose an encoder-based textual inversion method. During training, we fix the prompt as “a @ music with * as the rhythm”, where “@” and “*” respectively represent the placeholders for the genre and rhythm of the input dance. Our dual-path rhythm-genre inversion optimizes the rhythm encoder and genre encoder together to generate two text embeddings that correspond to placeholders, enabling control over the genre and rhythm of the generated music in a text-to-music generation pattern. The rhythm encoder integrates traditional feature processing techniques and projector networks. Parameters $v_{i}$, $v_{i@}$, and $v_{i}$ correspond to the text embeddings of the prompt, “@”, and “*”, respectively.
In this paper, we propose an encoder-based textual inversion technique, implemented as a separate network structure, that seamlessly integrates with text-to-music generation models. As shown in Fig. 2, we employ two popular text-to-music generation frameworks, namely the diffusion-based Riffusion [11] and the transformer-based MUSICGEN [8]. They are augmented with our dual-path rhythm-genre inversion, utilizing learnable pseudo-words that integrate rhythm and genre information to guide the dance-to-music generation. During training, the prompt remains fixed as “a @ music with * as the rhythm”, where “@” represents the genre description and “*” indicates the rhythm description. We obtain text embeddings for the two placeholder words using separate rhythm and genre encoders. By reconstructing the target audio, we simultaneously optimize the parameters of both encoders, facilitating their mutual enhancement. In the inference phase, flexible text descriptions can be used in addition to the fixed prompt to exert control over the high-level feature of music. The inclusion of the pseudo-word “*” representing rhythm is vital to ensure the generation of music that aligns with the dance.
3.1 Encoder-based Textual Inversion
Our objective is to leverage the generative capabilities of a pre-trained text-to-music model to generate music for dance videos. Typically, a text encoder firstly tokenizes a prompt into multiple indices, each corresponding to an embedding in the corresponding embedding lookup. These text embeddings are then encoded by a text transformer to serve as conditional guidance for the generative model. For Riffusion [11], the text encoder used is the text encoder of CLIP [34], while the generative model employed is the Latent Diffusion Models (LDMs). For MUSICGEN [8], the text encoder utilized is the T5 [35], while the generative model is the Transformer decoder.
One prevalent issue in most text-to-music models is their limited ability to control local attributes of music (e.g., rhythm), as these elements are challenging to describe accurately in natural language. This presents a significant challenge when generating music that aligns with dance videos, as precise control over the rhythm becomes crucial. Textual inversion is one of the personalization methods for image generation. It uses a pseudo-word (e.g., “*”), as a placeholder for a specific object and iteratively optimizes the text embedding associated with the pseudo-word. This process aims to integrate the textual description of the specific object into the vocabulary of the pre-trained text encoder, enabling targeted editing of the specific object. The optimization goal can be defined as follows:
[…]
where $z\sim E(x),\epsilon\sim\mathcal{N}(0,1)$, $\epsilon_{\theta}$ denotes LDMs, $c_{\theta}$ denotes the text encoder of CLIP and $y$ denotes the prompt.
It is natural to consider using a pseudo-word to represent the rhythm of a video and guide the generation of music, e.g., “lively melody with * as the rhythm and easy instrumental arrangement”. However, textual inversion requires training a separate model for each object, with fixed embeddings for all pseudo-words. Yet, the rhythm of dance is highly variable, and training a model for the rhythm of each video would be costly. Therefore, we propose an encoder-based textual inversion method, which allows us to define the rhythm as a variable attribute in inversion.
To enhance the expressiveness of individual pseudo-words, we introduce an encoder branch into the text encoder. This encoder extracts features from a specific input type (e.g., human poses), and maps them to the textual space. This enables a single pseudo-word with the capacity to exhibit distinct text embeddings contingent upon the input, allowing for control over the generation of diverse content. The encoder is iteratively optimized by reconstructing the target objects. For Riffusion, the loss function of the encoder is defined as follows:
$L_{E}=\mathbb{E}_{z,x,y,\epsilon,t}[\|\epsilon-\epsilon_{\theta}(z_{t},t,t_{\theta}(x,y))\|_{2}^{2}],$ (2)
where $x$ denotes a specific type of input, $t_{\theta}$ denotes the expanded text encoder. For MUSICGEN, the cross-entropy loss function of the encoder is defined as follows:
$L_{\text{ce}}=\frac{1}{K}\sum_{k=1}^{K}\text{CE}(G_{\theta}(t_{\theta}(x,y))_{k},{t}_{k}),$ (3)
where $t$ denotes target tokes, $x$ denotes a specific type of input, $t_{\theta}$ denotes the expanded text encoder, $G_{\theta}$ denotes the generative model and $K$ denotes the index of codebook.
With a single pseudo-word, it becomes feasible to exert precise control over specific attributes of the generated content (e.g., rhythm). This control can seamlessly coexist with human text editing, resulting in enhanced flexibility. The encoder-based textual inversion is not limited to a specific task but can be applied to different areas, where the attributes are variable and challenging to describe using natural language. In our present work, we demonstrate our approach by performing dual-path rhythm-genre inversion with rhythm and genre encoders to facilitate dance-to-music generation.
3.2 Rhythm Encoder
We combine traditional feature extraction with a Multilayer Perceptron (MLP). “Dance to the rhythm” serves as an important connection point for synchronizing music and video. Conversely, “drop the beat to the motion” is crucial for the task of creating music for dance videos. During dancing, dancers typically perform actions or transitions in sync with specific musical beats. The initiation and transitions of movements align with points of local maximum acceleration in kinematics. Building upon LORIS [51], we calculate the first-order difference of the 2D keypoints $p(t,j,c)$ of the dancers over time to obtain the motion velocities. We use $v_{x}$ and $v_{y}$ to represent the velocities in the x and y directions, respectively. Subsequently, we discretize them into $K$ intervals, a process which we refer to as direction discretization:
$v_{Q}(t,j,k)=Vl_{\theta}(t,j,k),$ (4)
$l_{\theta}(t,j,k)=\begin{cases}1,&\text{if }\text{$k=[\frac{\theta}{{2\pi}/{K}}]$}\\ 0,&\text{otherwise}\\ \end{cases}$ (5)
$V=\sqrt{{v_{x}}^{2}+{v_{y}}^{2}},\\ \theta=\arctan\left(\frac{v_{y}}{v_{x}}\right),$ (6)
where $\theta$ represents the direction angle, and $V$ represents the magnitude of the velocity. Continuing with the calculation of the first-order difference of $v_{Q}$, we obtain the discrete acceleration $a_{Q}$.
We retain the positive values of acceleration and then calculate the sum of the accelerations in the joint dimensions and directional dimensions to obtain the total acceleration, denoted as $a$:
$a=\sum_{j,k}a_{Q}(t,j,k).$ (7)
Now we determine the local maxima of $a$ within a time window. Then, we create a rhythm sequence $r$ where positions corresponding to the local maxima are set to 1, and all others are set to 0. We created a projector to map $r$ to the textual space, resulting in an embedding denoted as $v_{i*}$, which replaces the text embedding for the pseudo-word “*”.
3.3 Genre Encoder
Reconstructing the audio mel-spectrogram solely from rhythmic information may introduce music genre information into the rhythmic pseudo-word, which hinders attribute disentanglement. Furthermore, it also results in a lack of control in music generation based on dance genre. To overcome this, we introduce a genre encoder to enhance the diffusion model’s reconstruction capability by incorporating additional information. We represent dance genres using one-hot encoding. By employing a combination of linear and activation layers, we map the one-hot encoding to the textual space, effectively replacing the text embedding associated with the pseudo-word “@” with the corresponding genre embedding $v_{i@}$.
4 Experiments
Table 1: Qualitative comparison with state-of-the-arts methods. The best results are in bold and the second best results are underlined.
Rhythm Audio Quality Genre MOS
BCS $\uparrow$ BHS $\uparrow$ F1-score $\uparrow$ FAD $\downarrow$ Genre KLD $\downarrow$ Coherence $\uparrow$ Quality $\uparrow$
Ground Truth - - - - - 4.29 4.23
CMT [9] 0.3368 0.1515 0.2090 16.54 0.7297 1.81 3.25
CDCD [55] 0.4233 0.2151 0.2852 16.47 0.4592 1.96 1.41
LORIS [51] 0.2603 0.1073 0.1520 16.81 0.4405 1.66 1.33
MDM [45] 0.3798 0.4185 0.3982 4.812 0.5878 2.95 2.78
Ours (MUSICGEN) 0.4118 0.3874 0.3992 6.014 0.3354 3.52 3.59
Ours (Riffusion) 0.4761 0.4398 0.4572 3.416 0.3878 3.10 3.17
Figure 3: Qualitative examples of beat correspondence. The time scale interval is 0.1s. Our analysis of beats in real music and generated results from two frameworks show that the majority of the generated beats closely match the ground truth, with offsets below 0.2s. The Riffusion-based framework demonstrate slightly better beat alignment compared to MUSICGEN-based framework, consistent with quantitative metrics.
4.1 Experimental Setup
Dataset
In our experimental study, we utilize the primary perspective data from the AIST++ dance motion dataset [25]. This dataset comprises 70 songs paired with 460 Choreographies, evenly distributed across 10 different dance genres, along with corresponding 2D keypoints. We employ the official training, validation, and testing sets and segment the data into 5.12-second clips. As a result, our final training, validation, and testing sets consist of 2744, 36, and 36 samples, respectively. Furthermore, we transform the audio into corresponding mel-spectrograms to be used for experiments conducted on the Riffusion model. We utilize mmPose [7] to obtain 2D skeletons (dance motion sequences) of in-the-wild data from YouTube, using HRNet [44] pre-trained on the MS COCO [27] dataset.
Implementation details
We conduct experiments on our encoder-based textual inversion method using two prominent music generation frameworks: diffusion and transformer. The base models employed for these frameworks were Riffusion [11] and MUSICGEN [8], respectively. For Riffusion, we fix the parameters of the Latent Diffusion Models (LDMs) [36] and the text encoder, except for the encoder branches. We use the default hyperparameters of LDMs and employ a base learning rate of 0.0005. The model training is executed on an NVIDIA GeForce RTX 3090 GPU with a batch size of 1, taking approximately 12 hours to complete over 50 epochs. As for MUSICGEN, we fix the parameters of the audio compression model, transformer-based audio token decoder, and T5 encoder. We apply the AdamW optimizer with $\beta_{1}=0.9,\beta_{2}=0.95$, and a weight decay of 0.01. A warm-up learning rate of 1e-4 is implemented for all layers during the initial 6000 training iterations. Both encoders are jointly optimized. Training the model for 20 epochs using a single NVIDIA A40 GPU with a batch size of 1 requires approximately 10 hours.
4.2 Evaluation Metrics
Inspired by previous work and further analysis of the task, we develop a comprehensive evaluation protocol that incorporates multiple metrics to evaluate from the following perspectives.
Rhythm
Regarding the musical beats, a tolerance offset of 1 second is easily discernible to the human ear. To enhance the accuracy of evaluation, we refine the rhythm evaluation metrics by adjusting the tolerance offset from 1 second to 0.2 second. Specifically, we follow the setup of LORIS [51], where $B_{g}$ denotes the number of musical beats in the generated music, $B_{t}$ denotes the number of musical beats in the ground-truth music, and $B_{a}$ denotes the number of aligned musical beats. The evaluation metrics related to rhythm alignment are as follows:
  • Beat Coverage Score (BCS): the ratio of aligned musical beats to the generated musical beats ($B_{a}/B_{g}$);
  • Beat Hit Score (BHS): the ratio of aligned musical beats to the ground truth beats ($B_{a}/B_{t}$);
  • F1-score: an integrated assessment of rhythm alignment.
Audio quality
Frechet Audio Distance (FAD) [24] is widely used for evaluating audio quality by measuring the similarity between the generated audio and the ground truth. A lower FAD value indicates more close to the ground truth. We report the FAD based on VGGish audio embedding model [17], which is pre-trained on the YouTube-8M audio dataset [2].
Genre
We employ MS-SincResNet [6], a pre-trained genre classifier, to compute the genre Kullback Leibler Divergence (KLD) score. The KLD score measures the distance between the predicted genre probabilities of the generated music and the ground-truth genre labels, thereby evaluating the genre attributes of the generated music.
Table 2: Ablation study of our method. “CLAP” represents utilization of CLAP audio embeddings to replace genre labels. “MLP” represents the adoption of a Multi-Layer Perceptron (MLP) as the projector in the rhythm encoder. “Attn+Pos” denotes the employment of attention mechanism and positional encoding as the projector in the rhythm encoder.
Rhythm Audio Quality Genre
[…]
Ours (MUSICGEN) CLAP 0.3925 0.3155 0.3498 6.821 0.3593
MLP+Genre label 0.3958 0.3643 0.3794 6.469 0.3283
Attn+Pos 0.4118 0.3874 0.3992 6.014 0.3353
Ours (Riffusion) CLAP 0.4480 0.3085 0.3654 4.611 0.7053
MLP+Genre lable 0.4761 0.4398 0.4572 3.416 0.3878
Attn+Pos 0.3806 0.2768 0.3205 3.814  0.3821
4.3 Quantitative Evaluation
We conducted experiments by deploying our method in two mainstream frameworks: the diffusion-based Riffusion [11] and the transformer-based MUSICGEN [8]. Additionally, we compared our approach against four state-of-the-arts methods: CMT [9], CDCD  [55], LORIS [51], and MDM  [45]. Specifically, CMT is based on the transformer architecture, while the remaining methods utilize diffusion model. We re-implemented CDCD and LORIS using our dataset and evaluated CMT and MDM using the models provided by the original authors. Our experimental results based on two frameworks and their comparisons with other methods are shown in Table  1. Our experiments based on Riffusion achieve optimal results in terms of rhythm and audio quality while ranking slightly lower in genre performance. Experiments based on MUSICGEN demonstrate rhythm performance comparable to that of MDM and achieve optimal results in genre performance, but slightly weaker audio quality. The MIDI-based approach, CMT, performs significantly worse than our method across all metrics. Both CDCD and LORIS exhibit poor performance in terms of rhythm and audio quality. Specifically, CDCD demonstrates a decent BCS score but performs poorly in terms of BHS, indicating its inclination towards generating fewer yet accurate beats.
User study
To subjectively evaluate the performance of our method, we conduct a Mean Opinion Score (MOS) test. We provide participants with guidelines outlining the evaluation criteria for dance-to-music generation prior to the test. During the test, 78 participants rate a total of 70 generated samples, including 10 samples from each method and their corresponding ground truth. The evaluations are based on two criteria: dance-music coherence and audio quality. Dance-music coherence refers to the degree of alignment between the music and the dance video in terms of rhythm and style. Audio quality refers to the musicality of the audio (e.g., noise level, richness of melody). The rating scale ranges from 1 (best) to 5 (worst). As shown in Table 1, our approach based on MUSICGEN outperforms other methods in both metrics and achieves an acceptable margin compared to the ground truth. Interestingly, the results from MUSICGEN exhibit the best performance in dance-music coherence, indicating a tendency to generate music that closely align with human perception.
4.4 Qualitative Results
We visualized the beats of the generated music and compared them with the beats of ground truth, as shown in Fig. 3. Our method consistently generates reliable beats with offsets below 0.2, and the Riffusion framework shows superior beat alignment compared to MUSICGEN.
The generated audio samples, showcasing the comparison between our method and other approaches, can be accessed on the static webpage provided within the supplementary materials. The samples indicate that our model can generate more plausible dance music by guiding text-to-music models with prompts containing pseudo-words. Furthermore, our model demonstrates the capability to generate plausible beat-aligned dance music on in-the-wild data (e.g., film clips, virtual human dance videos, amateur dance videos). Furthermore, our method performs well in various other physical activities. Benefiting from the editability of the text-to-music model, our method exhibits the ability to generate diverse music that harmonizes with the dance video. The generated music aligns with the emotional atmosphere depicted in the videos, showcasing the robustness of our approach and its potential for real-world applications.
4.5 Ablation Study
Genre encoder
We first investigate the role of genre encoder that engage into dance-to-music generation. We introduce a variant called CLAP that replaces genre labels with audio embeddings derived from other segments of the same audio [50]. When the genre label is replaced with CLAP audio embeddings, experimental results indicate a decrease in all metrics. This observation indicates that the highly repetitive or monotonous audios in the AIST++ dataset is not suitable for facilitating the genre encoder in mapping CLAP audio embeddings to the textual space. As a result, achieving example-based genre control becomes challenging.
Rhythm encoder
We design two projectors to map rhythm sequences to the textual space: “MLP”, which utilizes a combination of multiple linear layers and activation layers as the projector, and “Attn+Pos”, which incorporates position embedding and attention modules as the projector. The results, as shown in Table  2, demonstrate that the Riffusion-based experiments perform better with the former projector and poorly with the latter. Conversely, the MUSICGEN-based experiments exhibit better performance with the latter projector. This indicates that MUSICGEN is more sensitive to position embedding.
4.6 Discussions and Limitations
Our approach generates diverse and high-quality music that aligns with videos, allowing for personalized, dynamic, and interactive music experiences in the context of dance videos. Furthermore, our method can be extended to various other physical activities (e.g., “jump rope”, “martial arts”, and “artistic gymnastics”). For limitation, our method currently supports only fixed-length video segments for music composition. The flexibility of accommodating variable-length segments would enhance the applicability of our method in real-world scenarios.
5 Conclusion
In this paper, we introduce an encoder-based textual inversion technique to seamlessly integrate rhythm and genre control into pre-trained text-to-music models. This approach offers a plug-and-play solution for text-to-music models. We further introduce an enhanced rhythm alignment evaluation metric and conduct a comprehensive assessment encompassing rhythm, audio quality, and genre. Experimental results demonstrate that our method can generate high-quality music that is synchronized with the beats across various genres, even with a limited dataset of 50 songs. Furthermore, our approach seamlessly adapts to in-the-wild data and enables flexible editing of high-level music attributes using text.
Future research may delve into the integration of multimodal motion information to achieve more robust beat generation results. Additionally, incorporating example-based genre control could provide more flexible user experiences in dance-to-music generation.